OpenStudy (anonymous):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -6 and 162, respectively.

OpenStudy (anonymous):

U there

OpenStudy (anonymous):

BRO

Miracrown (miracrown):

AFK

OpenStudy (anonymous):

Ok give me a sec just wanted to know if u were there

Miracrown (miracrown):

----- , 6 , ------ , ------, 162

Miracrown (miracrown):

So a geometric sequence is one determined by multiplying again and again by the same number. So we need to figure out what number to multiply -6 by two times and get 162 we could guess and check. Pick a number to start ...

OpenStudy (anonymous):

Sorry the lags men

Miracrown (miracrown):

Yea man, same here ;-;

OpenStudy (anonymous):

Take this as an example: I can't let you cheat

OpenStudy (anonymous):

Also, 2304/(-36) is the 5th term divided by the 2nd term which is -64 That is also: [(a)(r^4)] / [(a)(r)] = r^3 So, r^3 = -64 and r = -4 So, the 2nd term, -36, is (a)(r) = (a)(-4), so "a" = 9 We now have "a" and "r" and the general expression for the nth term is: (9)(-4)^(n-1) And remember that exponentiation takes precedence over multiplication, so only the "-4" is being raised to the "n-1" power.

OpenStudy (anonymous):

I know is a lot to throug at you.......